Teardown June 17 & 18

Teardown continues.  Yesterday (June 17) Clayton and Rich took down the clutter screen for the NCAS profiler at the Ops Center while Lou and Bill worked on the sounding site and packed equipment.  We all also loaded up the RASS surrounds onto the truck.   Emily (NCAS) arrived to last evening and today she, Clayton and Rich worked on the profiler, largely finishing up this afternoon, while Lou and Bill worked on the NCAR profiler site and continued loading up the seatainer.

Weather was been mainly clear and hot with light variable winds yesterday, however last night a storm came through with lots of lightning and some brief heavy rain.  Sadly about 30 miles to the northwest it was dry lightning which triggered a major wildfire with many lives lost.  Smoke and haze dominated the Perdigao area today.

Teardown June 16

Started teardown at the 1290 profiler site (ISS3).  Dismantled the profiler clutter screen and RASS, packed up the profiler electronics, the two computers at this site (profiler and data manager - backed up both before shutting down) and assorted communications, UPS, transformer and other components.  Also backed up and shutdown the sodar-RASS and it's computer (and backed that data up).  The sounding DM computer rebooted sometime today, probably because of overheating, so backed that computer up and shut it down.

The teardown crew is Lou, Clayton, Rich and Bill.

Weather very hot with no clouds, although some light breezes at times.  Some convection southeast of Castelo Branco in evening.

Daily Status June 15

Weather: hot with clear skies and light variable winds, picking up to a westerly breeze in the afternoon.  An extra sounding was launched last evening in support of TLS ops, plus an extra sounding this morning to aid the radiometer intercomparison at the Orange Grove site.  Surface met at sounding site still down. Todays soundings at 0Z, 6Z, 9Z, 12Z, and 18Z with the last sounding being in a Perdigao group photo.

The sodar went down again sometime this morning.  Was reset and running again around 14 UTC (see sodar blog entry comments).

The final day of the campaign!

Daily Status June 14

Weather: hot and clear with light variable winds.  One extra soundings this morning to SW wind obs of turbine wake, then another this evening plus regular soundings (0Z, 3Z, 6Z, 12Z, 18Z, 22Z).

Sodar went down again last night, restarted after noon sounding (see comment on sodar blog entry).  Gary reported that the sounding site DSM went down this afternoon - we all attempted to restart it, but no luck so the remaining 7 soundings (18Z on) will be without surface met data.

Kate left this morning and Gary this afternoon - thanks very much to both for all their hard work!


The sodar-RASS went down on June 12, apparently due to heat related issues.  A 220/110V transformer appears to have burnt out - since this runs the computer enclosures air conditioning, the computer overheated and shut-down.  In addition a temperature safety switch in in the antenna electronics box also tripped.  This is separate to the enclosure so shouldn't be related, but if the environment is hot enough they could both have failed on the same day.   Replaced the transformer with a temporary arrangement and reset the temperatue switch - the sodar was running again from about 16:30 UTC.

Daily Status June 13

Weather warm, light variable winds, gradually increasing scattered cumulus.  Regular soundings at 0Z, 6Z, 12Z, 18Z.   12Z sounding lost GPS at 2km.

Sodar plots not updating, will check on site later


Daily Status June 12

Weather hot with almost no clouds, and very light winds.  Regular soundings schedule of 0Z, 6Z, 12Z and 18Z.

Lou and Clayton arrived to begin preparing for teardown.

In the evening a storm developed near Castelo Branco and cloud increased.  Some outflow may have reached the Perdigao area.

Daily Status June 11

Warm, calm with scattered cirrus.  Extra soundings around morning transition, plus regular soundings (0Z, 3Z, 6Z, 9Z, 12Z, 18Z).  Missing data on some soundings due to unknown reasons, eg, 3Z lost at 400mb, 6Z had data gap 960-925mb, 12Z lost GPS between 800m and 6km.  Interestingly Notre Dame reported missing GPS from their iMet sondes this morning as well.

Vanda visited today, attending the 12Z sounding and doing a tour of various sites.  Afternoon science meeting after the regular briefing.

Daily Status 10 June

Weather: warm, mainly clear skies with a little thin cirrus, and no wind early but picking up to a light westerly in the afternoon.  Regular soundings at 0, 6, 12, and 18 UTC.  Additional soundings are planned for the Sunday morning transition.

Daily Status 9 June

Weather: warm with thin cirrus and light winds at surface.  Regular operations with soundings at 0, 6, 12, 18Z and extra launches at 15 ans 21Z.  A large group of locals (led by Jose Carlos and Anna) came to the 12Z sounding.

The 15Z and 18Z soundings both had problems.  For the 15Z sounding, Tyler (OU) reported a bad temperature sensor, and for the 18Z sounding Kate reported some kind of data lockup at 700mb - both soundings were relaunched.  Replaced desiccant in ground check unit after 2317Z sounding.

Isabel left this morning to return home - thanks very much Isabel (her first field project!).

Daily Status 8 June

Regular operations today with soundings at 0, 6, 12, 18Z.  Weather warm and light variable winds, thin upper level cloud.

The NCAS profiler is back up (see separate blog entry) after the trailer computer stalled.  The sounding DM computer rebooted this morning for unknown reasons, although the UPS was heard to beep briefly during the noon sounding so perhaps there were power fluctuations. 

Briefly paused 1290 Wind Profiler at 17UT to adjust RASS range since it's getting hot.  Shifted 5C warmer, so spect sig now 5C-35C, source 0C-40C.

The NCAS profiler shutdown for an undetermined reason around 20Z on June 6.   The trailer computer appeared to lock-up, but after a few reboot attempts came back up around 16Z on June 7.  It didn't appear to be an air conditioning or power issue, but Emily is connecting to attempt a diagnosis.  

Daily status 7 June 2017

4 good soundings by us today at 00, 06, 12, and 18 UTC, plus super soundings by OU at 03, 09, and 15 UTC, with one more to come at 21 UTC. Bill arrived today in time for the noon launch.

The NCAS profiler was down last night and today but is working again now.

Weather is hot and clear and mostly calm, looks like tomorrow will be just as warm.

Daily status 6 June 2017

4 good soundings today at 00, 06, 12, and 18 UTC. 

Super soundings tomorrow, so there will be launches every 3 hours from midnight on the 7th to midnight on the 8th.

Weather is warm and mostly clear with some breeze. 

Daily status 5 June 2017

4 good soundings today at 00, 06, 12, and 18 UTC.

Weather clear and wind from W/NW.